# Fork-Merge Decoding: Enhancing Multimodal Understanding in Audio-Visual Large Language Models

### v1

@@ 3.2 Decoding with Channel-wise Fusion in AV-LLMs
Masking strategy for input sequences. In channel-wise fusion models such as video-SALMONN sun2024video , the input sequence is structured as $\mathbf{x}=\{\mathbf{x}_{1}^{av},\ldots,\mathbf{x}_{U}^{av},\mathbf{x}_{U+1}^{l},\ldots,\mathbf{x}_{U+L}^{l}\},$ where $\mathbf{x}^{av}$ denotes audio-visual embeddings and $\mathbf{x}^{l}$ corresponds to language (prompt) embeddings. Here, $U$ and $L$ indicate the number of audio-visual and language sequence elements, respectively. To allow for modality-specific processing, we construct two masked variants of the input:
$\mathbf{x}^{v\text{[masked]}}=\mathbf{x}^{a\neg v}\oplus\mathbf{x}^{l},\quad\mathbf{x}^{a\text{[masked]}}=\mathbf{x}^{\neg av}\oplus\mathbf{x}^{l},$ (8)
where $\mathbf{x}^{a\neg v}$ and $\mathbf{x}^{\neg av}$ denote the video-masked and audio-masked embeddings, respectively, obtained by zeroing out the corresponding inputs before they are passed into the decoder layers.
Fork-merge decoding. $\mathbf{h}^{v}_{\text{fork}}$ and $\mathbf{h}^{a}_{\text{fork}}$ are obtained by passing the masked inputs $\mathbf{x}^{v\text{[masked]}}$ and $\mathbf{x}^{a\text{[masked]}}$ through the $L_{fork}$ layers (refer to Eq. (3)). To compute the merged hidden state $\mathbf{h}_{\text{merge}}$, we perform element-wise addition over the audio-visual embedding representations from both branches, based on the assumption that they capture complementary information. Since each branch processes modality-masked inputs (i.e., audio-masked for visual features and vice versa), their combination is expected to yield a more complete representation. Additionally, mean pooling is applied over the question prompt embedding positions to maintain consistency:
$\mathbf{h}_{\text{merge}}[i]=\begin{cases}\mathbf{h}^{v}_{\text{fork}}[i]+\mathbf{h}^{a}_{\text{fork}}[i],&\text{if }i\leq U,\\ \frac{1}{2}\left(\mathbf{h}^{v}_{\text{fork}}[i]+\mathbf{h}^{a}_{\text{fork}}[i]\right),&\text{if }i>U.\end{cases}$ (9)
We do not apply attention-guided fusion in the channel-wise setting, as the hidden states do not disentangle audio and visual embeddings. Finally, decoding is then continued by forwarding the merged hidden state $\mathbf{h}_{\text{merge}}$ through the remaining decoder layers to produce the final output logits (see Eq. (7)). Notably, FMD provides a unified processing framework that seamlessly integrates input masking, modality-specific processing via separate decoder branches, and merged decoding—making it broadly applicable to a wide range of AV-LLM architectures, regardless of their fusion strategies.
@@ 4 Experiments
We conduct extensive experiments to evaluate the effectiveness of the proposed Fork-Merge Decoding (FMD) method. This section begins with a description of the experimental setup, including baselines, datasets, and implementation details (Section 4.1). We then present both quantitative results (Section 4.2) and qualitative examples (Section 4.3) to demonstrate the effectiveness of our approach. Finally, we provide a deeper analysis of FMD in (Section 4.4).
@@
Model Decoding AVQA MUSIC-QA AVHBench
A→V V→A AV matching AV captioning
VideoLLaMA2 cheng2024videollama Vanilla 82.46$\pm$0.02 81.30$\pm$0.09 80.02 77.03 57.75 2.84$\pm$0.01
FMD 82.74$\pm$0.05 81.57$\pm$0.02 80.34 77.70 58.89 2.94$\pm$0.02
video-SALMONN sun2024video Vanilla 36.36$\pm$0.04 44.48$\pm$0.02 68.69 62.39 49.46 1.83$\pm$0.01
FMD 36.89$\pm$0.19 44.78$\pm$0.08 71.05 65.31 51.49 1.89$\pm$0.01
4.1 Experimental Setup
Baselines. We evaluate our approach using two representative AV-LLMs: VideoLLaMA2 cheng2024videollama and video-SALMONN sun2024video , which adopt token-wise and channel-wise fusion strategies, respectively. These models are chosen to demonstrate the generality of our proposed FMD method across different fusion paradigms. We initialize the models using their official pretrained weights.
Datasets and evaluation protocol. The AVQA yang2022avqa dataset contains 57,000 YouTube videos for evaluating real-world audio-visual understanding. MUSIC-AVQA li2022learning offers 45,867 QA pairs from 9,288 music performance videos, focusing on fine-grained audio-visual reasoning such as identifying sound sources and temporally aligning auditory and visual cues. AVHBench sung2024avhbench is the first benchmark specifically designed to assess audio-visual hallucinations in AV-LLMs. It comprises four subtasks: audio-driven video hallucination (A→V), video-driven audio hallucination (V→A), audio-visual matching (AV matching), and audio-visual captioning (AV captioning). For the transformer layer analysis in Section 4.4, we select 200 samples from each task (A→V, V→A, and AV matching), with the remaining data used for evaluation.
For the AVHBench dataset, the three binary (yes/no) tasks except AV captioning, we report classification accuracy. For AVQA, MUSIC-AVQA and AV Captioning, which involve open-ended responses, we follow the GPT-assisted evaluation protocol from the official VideoLLaMA2 implementation [^1] https://github.com/DAMO-NLP-SG/VideoLLaMA2/tree/audio_visual, reporting both the average score and standard deviation across multiple runs.
Implementation details. The number of layers used for fork phase is set to 5 for VideoLLaMA2 (out of 28 layers) and 8 for video-SALMONN (out of 40 layers). For attention-guided fusion, we set the weighting parameter to $\alpha=0.8$, which is a rounded value obtained from averaging fusion weights over 100 randomly sampled examples from AVHBench, as described in Section 3.1.
@@ 4.2 Quantitative Analysis
Comparison with vanilla decoding. To verify the effectiveness of our proposed FMD method, we evaluate it using VideoLLaMA2 and video-SALMONN on three datasets: AVQA, MUSIC-AVQA, and AVHBench. As shown in Table 1, FMD consistently outperforms the baseline across all tasks. This suggests that FMD enables the model to aggregate more balanced multimodal information. Especially, this is achieved by performing unimodal analysis within the fork layers and cross-modal understanding within the merge layers, allowing for more effective integration of visual and auditory inputs. Moreover, even though video-SALMONN is not trained on AVQA or MUSIC-AVQA, FMD still enhances performance, demonstrating the robustness and generalizability of our approach.
Table 2: Comparison of decoding methods on AVHBench sung2024avhbench dataset using VideoLLaMA2 cheng2024videollama . We compare the vanilla decoding, DoLa chuang2023dola VCD leng2024mitigating , SID huo2024self , and two FMD variants (Gaussian noise injection and zero-out masking). Among them, FMD with zero-out masking achieves the highest overall accuracy, underscoring its effectiveness in audio-visual reasoning.
@@
Vanilla - 80.02 77.03 57.75
DoLa chuang2023dola LLM 69.34 63.44 48.33
VCD leng2024mitigating VLM 75.96 69.67 52.52
SID huo2024self VLM 78.53 72.82 53.52
FMD w/ noise AV-LLM 79.06 78.23 57.10
FMD w/ zero-out AV-LLM 80.34 77.70 58.89
@@
Comparison with other decoding methods. We compare our FMD method with other test-time decoding strategies that operate at the logit level. The evaluation is conducted using the VideoLLaMA2 model on the AVHBench dataset. Specifically, the comparison includes the following: DoLa chuang2023dola , VCD leng2024mitigating , SID huo2024self , and FMD variants with Gaussian noise injection and zero-out masking. DoLa contrasts intermediate-layer outputs with final predictions to factually correct outputs. VCD reduces language bias by injecting Gaussian noise into the visual input and subtracting the resulting logits from the original ones. SID adopts a similar approach with VCD but preserves the least informative visual tokens based on attention maps, then contrasts their logits with the original outputs. For DoLa, we extract the intermediate layer’s output from the same layer specified in the original paper. For both VCD and SID, originally designed for vision-language models (VLMs), we adapt their methods to jointly handle audio and video modalities in AV-LLMs.
@@
As shown in Table 2, applying decoding strategies developed for LLMs or VLMs to AV-LLMs leads to degraded performance. This highlights the need for decoding methods that are specifically tailored to the unique characteristics of AV-LLMs, which differ substantially from those of unimodal or bimodal models. Additionally, injecting Gaussian noise into the inputs within FMD (denoted as FMD with noise) results in lower performance on A→V and AV matching tasks compared to vanilla decoding, although it does lead to improved performance on the V→A task. We attribute this to the inability of Gaussian noise to equally isolate modality-specific information, as further discussed in Supp. B. In contrast, zero-out masking within FMD yields consistent performance improvements across all tasks, demonstrating that the proposed FMD design is well-suited for audio-visual understanding.
@@
(c) An example of audio-visual matching (AV matching).
(d) An example of audio-visual captioning (AV captioning).
Figure 3: Qualitative results using VideoLLaMA2 cheng2024videollama on each task of the AVHBench sung2024avhbench dataset. FMD generates more accurate answers by effectively leveraging both audio and visual modalities, whereas vanilla decoding is more prone to confusion due to over-reliance on a single modality.
4.3 Qualitative analysis
Audio-driven video hallucination (Figure 3a). The model is asked a question based on video content, where a woman appears visually but is not the one speaking; instead, a separate female voice narrates the scene. Vanilla decoding wrongly assumes the visible woman is speaking, hallucinating audio-visual alignment. In contrast, FMD detects the mismatch and avoids hallucination.
Video-driven audio hallucination (Figure 3b). The video depicts a moving fire truck, but the audio contains only background chatter, with no siren or other relevant sounds. When asked whether the fire truck is making a sound, vanilla decoding incorrectly assumes the presence of sound based on visual priors. In contrast, FMD correctly identifies that the fire truck does not produce any sound.
Audio-visual matching (Figure 3c). This example presents a synthetically mismatched audio-video pair, where unrelated audio is overlaid on the video. Vanilla decoding fails to detect the inconsistency and incorrectly predicts a match. Unlike vanilla decoding, FMD better balances modality contributions and correctly identifies the mismatch between audio and visual inputs.
Audio-visual captioning (Figure 3d). This example requires a more complex understanding and interpretation of both video and audio, rather than a simple yes or no answer. Vanilla decoding relies on the video and merely infers that the saw is cutting a piece of wood. In contrast, when FMD is applied, it shows an understanding of the saw’s shape, the red line, and the cutting sound, indicating a deeper comprehension of the scene. This demonstrates the effectiveness of FMD in various audio-visual tasks. More qualitative results can be found in the Supp. C.
4.4 Further analysis
Figure 4: Layer-wise attention weight comparison on VideoLLaMA2 cheng2024videollama using 600 samples from the AVHBench sung2024avhbench dataset. We analyze the attention weights from the final token in the last decoder layer, focusing on the distribution across video and audio segments. Deeper merging within the network results in reduced attention to visual tokens and heightened attention to audio tokens.
@@ 4.4 Further analysis
Layer selection strategy for merge point. To determine the optimal layer for merging hidden states, we conduct two types of analysis. VideoLLaMA2 is used as the model, and 200 examples per task are randomly sampled from the AVHBench dataset across three tasks: A→V, V→A , and AV matching. To ensure a fair evaluation, all sampled examples are excluded from the test set,
First, we examine the attention weight distribution across layers to choose the merge point, as shown in Figure 4. We observe that the deeper the merge occurs in the network, the lower the attention paid to visual tokens and the higher the attention paid to audio tokens.
@@
Table 3: Ablation on fusion strategy. We ablate the fusion strategy in Eq. (4) with three methods: Exclusion, Average, and Attention-guided. Attention-guided fusion balances masked and unmasked inputs, yielding better performance.
Methods AVHBench sung2024avhbench
@@
Vanilla 80.02 77.03 57.75
Exclusion 80.02 77.79 57.10
Average 74.36 55.36 56.26
Attention-guided 80.34 77.70 58.89
Ablation on audio-visual fusion strategy. To demonstrate the effectiveness of our attention-guided fusion, we compare results with three different audio-visual fusion strategies in Table 3. We evaluate on the AVHBench dataset using the VideoLLaMA2 model.
In Eq. (4), the aggregation of masked and unmasked features is controlled by the fusion weight $\alpha$. Exclusion corresponds to setting $\alpha=1$, where the masked modality is entirely excluded. This leads to a performance drop compared to vanilla decoding. We attribute this to the causal nature of autoregressive models: fully ignoring one modality can disrupt the flow of information from previous tokens to the next prediction. Average fusion, where $\alpha=0.5$, also results in degraded performance—likely due to equal weighting between informative signals and noisy features. In contrast, when $\alpha$ is automatically determined via Eq. (5) and Eq. (6) by analyzing attention weights from the final decoder layer, it yields consistent performance gains across all tasks. This highlights the effectiveness of our attention-guided fusion strategy.
Table 4: Decoding speed comparison. FMD improves efficiency by using only the initial $L_{\text{fork}}$ layers in the fork phase.
Decoding Inference Speed $\downarrow$
(tokens/s)
Vanilla 0.34
VCD leng2024mitigating 0.69
DoLa chuang2023dola 0.54
FMD (Ours) 0.52
Decoding speed comparison. To assess the efficiency of our FMD method, we compare its decoding speed—measured in tokens generated per second—with other representative test-time decoding methods: VCD and DoLa. For this evaluation, we use VideoLLaMA2 and sample 100 examples from the AVHBench dataset.
As shown in Table 4, VCD exhibits the slowest decoding speed due to the requirement of an additional full model forward pass. DoLa is relatively faster, as it performs one full forward pass along with another partial forward pass up to an intermediate layer. FMD, on the other hand, requires two forward passes only up to the fork layers, after which the outputs are merged and passed through the remaining transformer layers. This makes FMD slightly faster than DoLa, while still achieving higher performance—as demonstrated in Table 2. These results suggest that FMD is designed with inference-time efficiency, and is also more effective than prior decoding methods originally developed for LLMs or VLMs.

### v2

@@ 3.2 Decoding with channel-wise fusion in AV-LLMs
Input masking strategy. In channel-wise fusion models such as video-SALMONN (Sun et al., 2024), the input sequence is structured as ${\bm{x}}=\{{\bm{x}}_{1}^{av},\ldots,{\bm{x}}_{U}^{av},{\bm{x}}_{U+1}^{l},\ldots,{\bm{x}}_{U+L}^{l}\},$ where ${\bm{x}}^{av}$ denotes audio-visual embeddings and ${\bm{x}}^{l}$ corresponds to language (prompt) embeddings. Here, $U$ and $L$ indicate the number of audio-visual and language sequence elements, respectively. To allow for modality-specific processing, we construct two masked variants of the input:
${\bm{x}}^{v\text{[masked]}}={\bm{x}}^{a\neg v}\oplus{\bm{x}}^{l},\quad{\bm{x}}^{a\text{[masked]}}={\bm{x}}^{\neg av}\oplus{\bm{x}}^{l},$ (5)
where ${\bm{x}}^{a\neg v}$ and ${\bm{x}}^{\neg av}$ denote the video-masked and audio-masked embeddings, respectively, obtained by zeroing out the corresponding inputs before they are passed into the decoder layers.
Fork-merge decoding. ${\bm{h}}^{\neg v}_{\text{fork}}$ and ${\bm{h}}^{\neg a}_{\text{fork}}$ are obtained by passing the masked inputs ${\bm{x}}^{v\text{[masked]}}$ and ${\bm{x}}^{a\text{[masked]}}$ through the $L_{fork}$ layers (refer to Eq. 2). To compute the merged hidden state ${\bm{h}}_{\text{merge}}$, we perform element-wise addition over the audio-visual embedding representations from both branches, based on the assumption that they capture complementary information. Since each branch processes modality-masked inputs (i.e., audio-masked for visual features and vice versa), their combination is expected to yield a more complete representation. Additionally, mean pooling is applied over the question prompt embedding positions to maintain consistency:
${\bm{h}}_{\text{merge}}[i]=\begin{cases}{\bm{h}}^{\neg v}_{\text{fork}}[i]+{\bm{h}}^{\neg a}_{\text{fork}}[i],&\text{if }i\leq U,\\ \frac{1}{2}\left({\bm{h}}^{\neg v}_{\text{fork}}[i]+{\bm{h}}^{\neg a}_{\text{fork}}[i]\right),&\text{if }i>U.\end{cases}$ (6)
We do not apply attention-guided fusion in the channel-wise setting, as the hidden states do not disentangle audio and visual embeddings. Finally, decoding is then continued by forwarding the merged hidden state ${\bm{h}}_{\text{merge}}$ through the remaining decoder layers to produce the final output logits. Notably, FMD provides a unified framework that seamlessly integrates input masking with modality-specific processing in separate decoder branches and merged decoding, making it broadly applicable to a wide range of AV-LLM architectures, regardless of their fusion strategies.
@@ 4 Experiments
4.1 Experimental setup
Baselines. We evaluate our approach using three representative AV-LLMs: VideoLLaMA2 (Cheng et al., 2024), video-SALMONN (Sun et al., 2024), and Qwen2.5-Omni (Xu et al., 2025).
Datasets and evaluation protocol. The AVQA (Yang et al., 2022) dataset contains 57,000 YouTube videos for evaluating real-world audio-visual understanding. MUSIC-AVQA (Li et al., 2022) offers 45,867 QA pairs from 9,288 music performance videos, focusing on fine-grained audio-visual reasoning such as identifying sound sources and temporally aligning auditory and visual cues. AVHBench (Sung-Bin et al., 2025) is the first benchmark specifically designed to assess audio-visual hallucinations in AV-LLMs. It comprises four subtasks: audio-driven video hallucination (A→V), video-driven audio hallucination (V→A), audio-visual matching (AV matching), and audio-visual captioning (AV captioning). For the transformer layer analysis in Section 4.4, we select 200 samples from each task (A→V, V→A, and AV matching), with the remaining data used for evaluation. For the AVHBench dataset, the three binary (yes/no) tasks except AV captioning, we report classification accuracy. For AVQA, MUSIC-AVQA and AV captioning, which involve open-ended responses, we follow the GPT-assisted evaluation protocol from the official VideoLLaMA2 implementation [^1] https://github.com/DAMO-NLP-SG/VideoLLaMA2/tree/audio_visual.
Implementation details. For attention-guided fusion, we set the weighting parameter $\alpha$ as described in Section 3.1, using 100 randomly sampled examples from the AVHBench dataset: $\alpha=0.8$ for VideoLLaMA2 and $\alpha=0.9$ for Qwen2.5-Omni. The number of layers used for the fork phase $L_{fork}$ is chosen to be roughly one-seventh of the total decoding layers: the 4th layer for VideoLLaMA2 and Qwen2.5-Omni (28 layers) and the 6th layer for video-SALMONN (40 layers). The rationale for these fork layer selections is further analyzed in Section 4.4.
Table 1: Comparison of audio-visual understanding performance. We evaluate FMD on VideoLLaMA2, video-SALMONN, and Qwen2.5-Omni. Vanilla denotes the original decoding strategy of each model. Experiments are conducted on AVHBench, AVQA, and MUSIC-AVQA datasets. FMD consistently improves performance across all benchmarks, especially in AV matching and captioning. For Qwen2.5-Omni, we leave the MUSIC-AVQA entry as N/A due to out-of-memory issues.
Model Decoding AVHBench AVQA MUSIC-AVQA
@@ 4.1 Experimental setup
VideoLLaMA2 Vanilla 80.02 77.03 57.75 2.84 60.23 81.30
FMD 80.45 77.52 59.01 2.95 61.46 81.50
video-SALMONN Vanilla 68.69 62.39 49.46 1.83 28.20 44.48
FMD 70.51 65.41 54.77 2.01 40.46 50.60
Qwen2.5-Omni Vanilla 80.77 71.20 77.45 3.25 86.49 N/A
FMD 81.30 71.77 78.22 3.36 86.61 N/A
@@ 4.2 Quantitative analysis
Comparison with vanilla decoding. To verify the effectiveness of our proposed FMD method, we evaluate it with VideoLLaMA2, video-SALMONN, and Qwen2.5-Omni on three datasets: AVQA, MUSIC-AVQA, and AVHBench. As shown in Table 1, applying FMD consistently improves performance over vanilla decoding, which represents the original model inference, across all tasks. Notably, the gains are more pronounced in the AV captioning task (evaluated on a 5-point scale), which requires generating long and sophisticated answers, with relative improvements ranging from 3.4% to 9.8% across models. Among the models, video-SALMONN benefits the most, achieving a remarkable 12.26% increase on the AVQA dataset. This is particularly significant because video-SALMONN has not been trained on AVQA or MUSIC-AVQA, yet FMD still enhances its zero-shot inference. These results highlight the robustness and strong generalizability of our approach.
Table 2: Comparison of decoding methods on AVHBench dataset using VideoLLaMA2. We compare vanilla decoding with DoLa, VCD, SID and two FMD variants (Gaussian noise injection and zero-out masking). Among them, FMD with zero-out masking achieves the highest overall accuracy, underscoring its effectiveness.
@@
[…]
DoLa (Chuang et al., 2024) LLM 69.34 63.44 48.33
VCD (Leng et al., 2024b) VLM 75.96 69.67 52.52
SID (Huo et al., 2025) VLM 78.53 72.82 53.52
FMD w/ noise AV-LLM 79.17 78.03 57.76
FMD w/ zero-out AV-LLM 80.45 77.52 59.01
@@
Comparison with other decoding methods. We compare our FMD method with other test-time decoding strategies that operate at the logit level. The evaluation is conducted using the VideoLLaMA2 model on the AVHBench dataset. Specifically, the comparison includes the following methods: DoLa (Chuang et al., 2024), VCD (Leng et al., 2024b), SID (Huo et al., 2025), and FMD variants with Gaussian noise injection and zero-out masking. DoLa contrasts intermediate-layer outputs with final predictions to factually correct outputs. VCD reduces language bias by injecting Gaussian noise into the visual input and subtracting the resulting logits from the original ones. SID adopts a similar approach with VCD but preserves the least informative visual tokens based on attention weights, contrasting their logits with those of the original outputs. For DoLa, we follow the original setup and extract outputs from the same intermediate layer reported in the paper. For VCD and SID, which were originally developed for vision-language models (VLMs), we adapt their procedures to handle both audio and video modalities in AV-LLMs.
@@
As shown in Table 2, applying decoding strategies developed for LLMs or VLMs to AV-LLMs leads to degraded performance. This highlights the need for decoding methods that are specifically tailored to the unique characteristics of AV-LLMs, which differ from those of unimodal or bimodal models. Additionally, injecting Gaussian noise into the inputs within FMD (denoted as FMD with noise) results in lower performance on A→V and AV matching tasks compared to vanilla decoding, although it does lead to improved performance on the V→A task. We attribute this to the inability of Gaussian noise to equally isolate modality-specific information, as further discussed in Supp. A.3. In contrast, zero-out masking within FMD yields consistent performance improvements across all tasks, demonstrating that the proposed FMD design is suitable for audio-visual understanding.
(a) An example of audio-visual matching from AVHBench.
(b) An example of audio-visual question & answering from AVQA.
Figure 3: Qualitative results with VideoLLaMA2 on AVHBench and AVQA. Vanilla decoding often relies on a single modality, resulting in incomplete or inconsistent outputs, whereas FMD effectively integrates both audio and visual information to produce more accurate and coherent results.
4.3 Qualitative analysis
Audio-visual matching from AVHBench (Figure 3a). The video shows a man fishing in a stream, and the audio contains the sound of water flowing in a valley with occasional splashes from footsteps. The vanilla decoding mainly captures the visual content, describing only the presence of a man fishing in a stream. However, it overlooks the acoustic context, failing to reflect the sound of flowing water and splashes. By contrast, our proposed FMD generates a more comprehensive description that integrates both modalities, enriching the visual detail (e.g., specifying the fishing rod) and the audio detail (e.g., capturing the sounds of water flow and splashes). This highlights the strength of FMD in balancing and fusing multimodal cues, leading to richer and more faithful outputs.
Audio-visual question & answering from AVQA (Figure 3b). The video shows a train, but in the early frames its appearance could be confused with a cable car. The accompanying audio contains the sound of a train gradually accelerating as it departs from the station. The vanilla model, relying mainly on the ambiguous visual information, incorrectly answers (A) Cable car. In contrast, FMD correctly identifies the scene as a (C) Train, as it integrates both visual and auditory cues. This illustrates that FMD effectively leverages audio information to resolve visual ambiguity, resulting in more accurate and context-aware answers. More qualitative results, highlighting the ability of FMD to capture both audio and visual information, can be provided in the Supp. B.
4.4 Further analysis
Figure 4: Layer-wise hidden state similarity in VideoLLaMA2. $L_{\text{fork}}$ is chosen from the early stage.
Layer selection for merge point. To determine the optimal layer for merging hidden states, we first measure the similarity across all hidden states of VideoLLaMA2, following the approach proposed in (Sun et al., 2025b). Their study on LLMs shows that layers can typically be organized into 4–5 clusters: 1–2 clusters in the early stage, a large cluster in the middle, and 1–2 clusters in the later stage. In the early layers, modalities align in feature space, with intra-modal encoding strengthened and inter-modal interaction suppressed. In the middle layers, deeper modality fusion occurs, while in the later layers, it prepares task-specific outputs. As shown in Figure 4, we observe a similar clustering pattern in VideoLLaMA2. Based on this observation, we select the fork layer $L_{fork}$ from the early stage, where feature alignment and intra-modal encoding occur (Wei et al., 2024; Yu & Lee, 2025). This choice matches our decoding strategy, aiming to strengthen the independent intra-modal representations of vision and audio.
[…]
Table 3: Ablation of fusion strategies in Eq. 3 with VideoLLaMA2. Attention-guided fusion (Fixed) balances masked and unmasked inputs, improving performance.
Methods AVHBench
@@ 4.4 Further analysis
Exclusion 79.38 76.79 54.83
Average 75.11 55.31 57.16
Attention (Adaptive) 79.49 72.78 59.19
Attntion (Fixed) 80.45 77.52 59.01
Ablation on audio-visual fusion strategy. To demonstrate the effectiveness of our attention-guided fusion, we compare it with three alternative audio-visual fusion strategies on the AVHBench dataset using the VideoLLaMA2 model in Table 3.
@@
Figure 5: Layer-wise attention weight comparison on VideoLLaMA2 using 600 samples from the AVHBench dataset. We analyze the attention weights from the final token in the last decoder layer, focusing on the distribution across video and audio segments. Deeper merging within the network results in reduced attention to visual tokens and heightened attention to audio tokens.
@@
Table 4: Decoding speed comparison. FMD achieves faster inference since the process after $L_{\text{fork}}$ matches the original model.
@@
(sec/token)
[…]
SID 0.71 2.09
VCD 0.69 2.03
DoLa 0.54 1.59
FMD (Ours) 0.43 1.26
Decoding speed comparison. To validate the efficiency of FMD, we compare its inference speed against three representative test-time decoding methods, SID (Huo et al., 2025), VCD (Leng et al., 2024b) and DoLA (Chuang et al., 2024). We apply each decoding method to VideoLLaMA2 and measure the time required to generate a single token, reporting both the absolute latency in seconds per token and the relative speed normalized to vanilla decoding in Table 4. The results are obtained on 100 examples from the AVHBench dataset.
SID and VCD exhibit relatively slow decoding speed. Since they require two full forward passes, one on the original input and one on the modality-corrupted input to contrast the resulting logits, the computational cost nearly doubles. DoLA achieves faster inference than SID and VCD by leveraging intermediate layer outputs to refine the model’s predictions, thereby mitigating the inefficiency of repeated full forward passes. Our proposed FMD achieves the fastest inference speed among the three methods. This efficiency stems from the fact that only the fork layers require dual forward passes, making FMD not only computationally efficient but also more effective than prior decoding methods originally developed for LLMs or VLMs, as evidenced by its superior performance reported in Table 2. Additional experiments analyzing how inference speed varies with deeper $L_{fork}$ settings are provided in Supp. A.4.